Hockey team receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 58 points. They won 11 more games than they tied. How many wins and how many ties did the team have?
Answer #1
win=w
tie=t
They won 11 more games than they tied
w=11+t
receive 2 points when they win and 1 point when they tie.
2*w+1*t=58
2*(11+t)+t=58
3*t=58-22=36
t=12
w=11+t=23